OpenStudy (amonoconnor):

I'm stuck on a problem dealing with "e", and using Implicit Differentiation to find dy/dx. I've tried solving this one, but am not sure what the correct process is, if I'm not doing it right.. ? The problem: e^(x/y) = x-y; Find dy/dx with Implicit Differentiation. Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

@AryaVegapunk :)

zepdrix (zepdrix):

You remember your (e^x)' derivative? :)

zepdrix (zepdrix):

\[\large\rm e^{x/y}=x-y\]

OpenStudy (amonoconnor):

it's just e^x, right?

zepdrix (zepdrix):

Good. So same thing with e^(x/y), but apply chain rule:\[\large\rm \frac{d}{dx}e^{stuff}=e^{stuff}(stuff)'\]

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}e^{x/y}=e^{x/y}\left(\frac{x}{y}\right)'\]You could apply quotient rule maybe, unless you prefer product rule.

OpenStudy (amonoconnor):

So... \[(e^{\frac{x}{y}})*[(1)(y^(-1) + x(-y^(-2)dy/dx] = (1-1(dy/dx))\]

OpenStudy (amonoconnor):

Lols, I prefer Product Rule. And my laptop just died so I'm in my iPod touch, just FYI if I'm slower now.

zepdrix (zepdrix):

\[\large\rm e^{x/y}\left(x\cdot y^{-1}\right)'=e^{x/y}\left[1\cdot y^{-1}-1x\cdot y^{-2}y'\right]\]So you went with product rule? Ok cool. Hmm I'm not sure about your last simplification though..

zepdrix (zepdrix):

Just a note, when you do exponents in the equation tool, you have to do ^{ } not ^( )

OpenStudy (amonoconnor):

Is that right so far?

zepdrix (zepdrix):

\[\large\rm e^{x/y}=x-y\]\[\large\rm e^{x/y}\left[y^{-1}-xy^{-2}y'\right]=1-y'\]Oh oh you were differentiating the right side, ok great, looks good so far

OpenStudy (amonoconnor):

The derivative of the right side, (x-y)? That's not (1-(dy/dx))?

zepdrix (zepdrix):

oh boy you are slow XD haha np

OpenStudy (amonoconnor):

Gotcha! Thanks!

zepdrix (zepdrix):

Now you need to get all of your y' to the same side.

OpenStudy (amonoconnor):

Yeah, the app refreshes like molasses..

zepdrix (zepdrix):

Distribute the exponential,\[\large\rm e^{x/y}y^{-1}-e^{x/y}xy^{-2}y'=1-y'\]The one with the y', add it to the other side, and also subtract 1 to the left side, \[\large\rm e^{x/y}y^{-1}-1=e^{x/y}xy^{-2}y'-y'\]Factor our your y'\[\large\rm e^{x/y}y^{-1}-1=(e^{x/y}xy^{-2}-1)y'\]Then divide, ya? :d

zepdrix (zepdrix):

Quotient rule woulda been a little nicer I think. These negative exponents are so ugly >.<

OpenStudy (amonoconnor):

e^(x/y)*y^(-1) - e^(x/y)*xy^(-2)(dy/dx) = 1 - dy/dx

OpenStudy (amonoconnor):

Can I do anything more than this? :/